# Enhancing Systematic Decompositional Natural Language Inference Using Informal Logic

### v2

@@ Abstract
Contemporary language models enable new opportunities for structured reasoning with text, such as the construction and evaluation of intuitive, proof-like textual entailment trees without relying on brittle formal logic Tafjord et al. (2022); Weir et al. (2023). However, progress in this direction has been hampered by a long-standing lack of a clear protocol for determining what valid compositional entailment is. This absence causes noisy datasets and limited performance gains by modern neuro-symbolic engines. To address these problems, we formulate a consistent and theoretically grounded approach to annotating decompositional entailment datasets, and evaluate its impact on LLM-based textual inference. We find that our resulting dataset, RDTE (Recognizing Decompositional Textual Entailment), has a substantially higher internal consistency (+9%) than prior decompositional entailment datasets, suggesting that RDTE is a significant step forward in the long-standing problem of forming a clear protocol for discerning entailment. We also find that training an RDTE-oriented entailment classifier via knowledge distillation and employing it in a modern neuro-symbolic reasoning engine significantly improves results (both accuracy and proof quality) over other entailment classifier baselines, illustrating the practical benefit of this advance for textual inference.
Figure 1: (Upper) Two hypothesis decompositions suggested by an LLM. The first makes an argument that is generally acceptable to a human. The second contains a fact that is not always true and another that is irrelevant to the entailment. Recognizing such an invalid decomposition is core to recent neuro-symbolic reasoning algorithms, but LLMs struggle at the task. (Lower) Ambiguous definitions of entailment have hampered progress in annotating data to improve the models. (b) We find that a faceted definition yields both a clean dataset (RDTE) and significant downstream performance improvements.
@@ 1 Introduction
What denotes a deductively valid explanation for an inference? While formal logicians might support a stringent condition of completeness according to well-defined formal axioms, evidence suggests that humans do not think of explanation validity so stringently Sulik et al. (2021), accepting incomplete explanations that “seldom capture the complete deductive processes from a set of axioms to a statement” Tan (2022). Recognizing textual entailment (RTE; Dagan et al., 2005), which generally centers around emulating a human’s opinion of entailment, commits not to a single definition of validity but rather collects human preferences on a wide spectrum of criteria from “strict deductive inferences” to “very implicit and fallible common-sense inferences” Gubelmann et al. (2023).
The lack of a well-defined task definition creates a challenge for recent lines of work on explainable AI, much of which seeks to make the reasoning processes of complex systems like large language models (LLMs) transparent and convincing to users. One such line of work is on entailment tree-based reasoning Dalvi et al. (2021); Bostrom et al. (2022); Tafjord et al. (2022). Most entailment tree-based literature has avoided the critical question of reasoning validity: whether each step of a model-generated tree is a valid argument to a human in support of its hypothesis. Existing evaluations generally focus on tree reconstruction metrics Ribeiro et al. (2023a) or end-task QA performance, and do not consider whether the quality of model-generated trees necessarily aligns with the accuracy of the answers. It is entirely possible—and indeed, not uncommon—for a system to arrive at the correct conclusion while presenting a flawed decomposition. This disconnect motivates our investigation into the automatic detection of errors within proof trees. Reliably performing decompositional RTE undergirds a system’s capacity to (1) be right for the right reasons and (2) not be wrong for the wrong reasons.
@@
The existing discourse surrounding a proper definition of NLI (e.g. Manning (2006)) has thus far not touched on specifically decompositional entailment, particularly governing the sorts of decompositions encountered during a recursive entailment tree search algorithm. Towards addressing this lack of clarity, we propose an evaluation centered around the notion that a valid decomposition is a valid argument for why the hypothesis should be believed. We take inspiration from the “Relevance, Acceptability, and Sufficiency” criteria for a logically good argument developed in the field of informal logic Johnson and Blair (1977); Groarke (2022), and design a novel approach to reasoning about entailment in a principled manner. We release this protocol along with a collection of over 1000 associated expert annotations. Through experiments, we find that nearly all models trained on previous compositional entailment datasets and LLMs like GPT fall short of human-level performance on our challenge set, which we term RDTE (Recognizing Decompositional Textual Entailment).
We find that under our prompting protocol, GPT-4 can serve as a “teacher” in a novel knowledge distillation pipeline that annotates the reasoning traces of a non-optimized entailment reasoner in a given domain. We collect and release a large artifact (24K items per domain) of GPT-4’s annotations over these traces for use by future work.
@@
We illustrate the effectiveness of student models trained via this pipeline as the linchpin of TreeWise, an entailment tree engine inspired by Nellie but which supports in-context learning, forward chaining, branch consolidation, and most importantly, improved decompositional entailment recognition. TreeWise not only surpasses entailment tree-producing approaches on established benchmarks like EntailmentBankQA, but also adapts to complex tasks such as HotpotQA that require reasoning over less structured knowledge sources such as Wikipedia. We show that using the knowledge-distilled student model improves TreeWise’s QA task performance, and raises the overall quality of the entailment trees produced by the system. Our contributions are therefore:
@@
  • A new entailment challenge set, RDTE, crafted via an informal logic-inspired protocol, that tasks models to verify the validity of a hypothesis decomposition
@@ 2.1 Background: RAS Criteria
A seminal framework developed by informal logicians to replace strict deductive logic criteria is known as RAS: Relevance, Acceptability, and Sufficiency Johnson and Blair (1977). Noting that each element is subject to substantial academic debate (different systems of informal logic define them differently), we review each criterion as we interpret them for this work:
Relevance of premise A concerning conclusion B is defined as whether the truth of A makes a difference to the truth of B. Blair (2012). The extent of this difference can vary; e.g. whether (A) “the earth has oxygen” is true technically has relevance to (B) birds can fly (since birds breathe oxygen), but is less relevant than (A’) birds have wings. Relevance is of particular interest to a recursive reasoner, as one would not want to waste time proving an irrelevant statement like (A) when trying to compositionally prove (B).
@@
Sufficiency is “the property of an argument’s premises of supplying all the grounds that are needed to make it reasonable to believe its conclusion.” Johnson and Blair (1977). This is left intentionally vague; Blair (2012) admits “the criterion of sufficiency, for justificatory arguments, is best seen as a placeholder for whatever version and standards of sufficiency are appropriate for the particular situation in question.” In this way, we should consider sufficiency to be a question and problem-specific criterion. This suggests the key observation that the grounds for valid entailment are inherently domain-specific. Blair (2012) also notes the dependent relationship between sufficiency and the other two criteria: that a sufficient argument presumes acceptability and relevance.
@@ 2.2 Implementing RAS for RTE Annotation
We draw inspiration from these criteria to construct a protocol that investigates a decomposition like the ones shown in Figure 1 for each SAR component in turn. An important aspect of these normative terms is that they are all scalar: a premise can be more or less relevant and acceptable, and the sufficiency of an argument composed of premises could always be strengthened by adding more premises. In a departure from works such as Tafjord et al. (2022) who collect binary factuality and “reasoning correctness” judgments, we collect RAS judgments on an ordinal scale Zhang et al. (2017) from 1 to 5, where each score is assigned a specific set of conditions. We also collect an ordinal judgment for redundancy, which is not strictly a component of RAS assessment; we observe that redundant premises are particularly problematic for entailment tree search, as proving the same information twice wastes search budget. We implement redundancy as “conditional irrelevance”:
@@
The sufficiency label, which most directly resembles that of a typical RTE label, is dependent on the first two following Blair (2012)’s observation. Exact directions can be found in §B, including a substantial list of conditions around a decomposition that indicate what the score should be.
@@ 3 Data Collection
3.1 Generating Decompositions
We seek to construct a dataset of decompositional entailment judgments that are of the kind that a reasoning system might come across while performing QA. Following Tafjord et al. (2022), we found it apt to annotate decompositions generated during a model’s reasoning search traces over training questions. We use the backward chaining reasoning process of TreeWise, which is introduced in §5. Each item is a hypothesis and a set of 2-3 premises that the model proposes might conjunctively entail the hypothesis in the context of a given question. To collect a representative sample of hypotheses, we pull from three classes:
@@ 3.1 Generating Decompositions
    Top-level correct hypotheses representing the right answer options for multiple-choice questions. These are important to annotate so that the model is right for the right reasons.
@@
    Recursive correct hypotheses generated by an LLM as premise subqueries for the top-level correct hypotheses
@@
    Top-level incorrect hypotheses representing the incorrect answer option deemed by GPT-4 to be closest to correct. These are important to annotate so that the model is not wrong for the wrong reasons.
@@
Annotation Process
Four of this paper’s authors, all highly proficient or native English speakers, annotated 1000 decompositions total from the two domains with two-way redundancy. One author was an annotator on all items. We initially annotated a handful of examples together to develop a list of around 30 conditions to check for in a decomposition that indicate certain RAS scores. We then annotated the rest of the dataset independently.
While the dataset is annotated for sufficiency on a 5-point scale, we suggest using a threshold of $\geq 4$ if one wants to evaluate models under binary entailment metrics. This corresponds to items for which the only acceptable flaws are (A) a 3rd redundant premise given an otherwise perfect 2-premise entailment or (B) minor missing/implicit information that does not affect layman reasoning. To arrive at a single clean entailment label for evaluation, we reconciled disagreements by discussing all items for which its two annotated sufficiency scores were on either side of 3.5. Anecdotally, we found that a vast majority disagreements were due to human error, e.g. missing a condition or not recognizing a particular flaw in reasoning.
Comparison to Existing Data
Prior to reconciling disagreements, we measured raw annotation agreement and compared it to recent attempts to collect decompositional entailment labels. Tafjord et al. (2022) collected annotations for 3.7K decompositions generated by Entailer, while Clark et al. (2023) collected annotations for 24.4K items generated by GPT3.444The BaRDa dataset comprises a specifically filtered subset of these two datasets that exhibit maximal annotator agreement. We obtained the pre-filtered data from the authors for the purposes of comparison with our work. The instructions given to annotators for these datasets are very high-level, simply asking whether the “reasoning goes wrong.” We believe this creates a much lower threshold for acceptability than the RDTE protocol; these datasets are majority labeled “entailment,” but on manual inspection we found numerous items that exhibited redundant, irrelevant, and fallacious arguments.
Nevertheless, we found RDTE to have a higher internal annotator consistency rate than these datasets, with a rate of 79% compared to 70% for the GPT3 and 61% for the Entailer data.
3.2 RDTE Analysis
@@
P1: Frank Vincent appeared in The 2005 film Remedy.
5 5
P2: Frank Vincent appeared in The Sopranos and several mob movies
5 5
P3: Ralph De Vito directed several mob movies.
5 5
@@
P1: Changes in the Grand Canyon’s landscape include becoming wider.
5 5 5
P2: Erosion is one of the processes that can change a landscape.
5 5 5
P3: The Grand Canyon is a landscape.
5 3 5
@@
P1: Ordos City is predominantly rural.
- 3 5
P2: Predominantly rural areas in China are often found in the western part of the country.
- 3 5
@@
Table 1: Example RDTE annotations.
@@ 4 RDTE Evaluation
We elicit itemwise decomposition quality judgments from a series of methods based on the RDTE protocol as well as existing methods for judging decompositional entailment. For methods that provide factuality judgments as part of their inference process, we turn off this functionality for the Hotpot subset of RDTE (this entails either rewriting a prompt or not making a second model call).
GPT Methods
We test the following prompt-based methods, using both GPT-4 and ChatGPT.555gpt-4-0613 and gpt-3.5-turbo-1106 All prompts can be found in the appendix.
  • An ICL prompt containing the RDTE annotation rubric and 4 example batches (10-15 decompositions per hypothesis) and a Zero-Shot prompt containing only the RDTE rubric
  • The prompts used by Clark et al. (2023) to evaluate models on the BaRDa dataset. These are separate prompts for the entailment and premise factuality judgments.
Existing Methods
We test the T5-based Entailer-11B model from Tafjord et al. (2022), which is trained in a multi-angle fashion to alternatively provide entailment and factuality judgments. We also test the similar Nellie-3B model, as well as the fine-tuned RoBERTa classifier used by Nellie as a secondary entailment filter. We also take an off-the-shelf (OTS) RoBERTa finetuned for NLI using a large suite of 600 datasets [^6] https://huggingface.co/sileod/deberta-v3-large-tasksource-nli that includes QASC Khot et al. (2020) and ARC.
Knowledge Distillation
Search algorithms like Nellie make hundreds of calls to NLI models for every question, making large, slow models like GPT-4 unrealistic for use as an entailment filter. We instead test whether we can use knowledge distillation to train smaller student models to imitate GPT-4’s performance. We extracted 20K RDTE judgments from GPT-4 over 2.3K hypotheses in each of the ARC and Hotpot domains. We then used this silver data to fine-tune (A) the base RoBERTa model finetuned to serve as Nellie’s filter using the Entailer and QASC datasets, and (B) ChatGPT using OpenAI’s fine-tuning API. To promote token efficiency, the GPT-4 teacher and ChatGPT student perform classification over a batch of decompositions for a single hypothesis in each prompt. See Table 4 for how batching marginally affects GPT-4’s performance on RDTE.
@@
ChatGPT (RDTE Zero-Shot) 40 ${}^{\dagger}$
ChatGPT (BaRDa) ${}^{\dagger}$ 34
@@
ChatGPT 48 (${}^{\dagger}$ ${}^{\dagger}$
RoBERTa 66 (${}^{\ddagger}$ ${}^{\ddagger}$
[…]

### v3

@@ Abstract
Recent language models enable new opportunities for structured reasoning with text, such as the construction of intuitive, proof-like textual entailment trees without relying on brittle formal logic Tafjord et al. (2022); Weir et al. (2023). However, progress in this direction has been hampered by a long-standing lack of a clear protocol for determining what valid compositional entailment is. This absence causes noisy datasets and limited performance gains by modern neuro-symbolic engines. To address these problems, we formulate a consistent and theoretically grounded approach to annotating decompositional entailment and evaluate its impact on LLM-based textual inference. We find that our new dataset, RDTE (Recognizing Decompositional Textual Entailment), has a substantially higher internal consistency (+9%) than prior decompositional entailment datasets. We also find that training an RDTE-oriented entailment classifier via knowledge distillation and employing it in an entailment tree reasoning engine significantly improves both accuracy and proof quality, illustrating the practical benefit of this advance for textual inference.
Figure 1: (Upper) Two hypothesis decompositions suggested by an LLM. The first makes an argument that is generally acceptable to a human. The second contains a fact that is not always true and another that is irrelevant to the entailment. Recognizing such an invalid decomposition is core to recent neuro-symbolic reasoning algorithms, but LLMs struggle at the task. (Lower) Ambiguous definitions of entailment have hampered progress in annotating data to improve the models. We find that a faceted definition yields both a clean dataset (RDTE) and significant downstream task improvements.
@@ 1 Introduction
Textual inference using entailment trees has emerged as a promising avenue for explainable AI in domains requiring complex reasoning Dalvi et al. (2021); Bostrom et al. (2022). Inspired by classical symbolic reasoners that produce proof trees of a hypothesis, recent entailment tree systems (e.g. Nellie Weir et al. (2023) or IRGR Ribeiro et al. (2022)) explain a hypothesis by constructing, via an LM-based generative search algorithm, a tree of recursive natural language (NL) decompositions whose children entail their parents. Entailment trees provide a natural and granular medium for explaining complex multi-step decisions to users and create opportunities to ground reasoning to large NL knowledge corpora Weir et al. (2023).
Entailment tree reasoners are most useful to humans if they are right for the right reasons: They not only give correct answers, but each tree step is a valid deduction supporting its hypothesis. While some tree generators leverage classifiers for recognizing textual entailment (RTE; Dagan et al., 2005) to identify which hypothesis decompositions are illogical and should be discarded Yang et al. (2022), it is still common for systems to arrive at the correct conclusion while presenting a flawed decomposition, or to accept incorrect conclusions based on faulty logic. Thus, the task of recognizing decompositional textual entailment is a major barrier to progress on tree-based reasoners.
Work on entailment trees has largely avoided this critical question of reasoning validity. Existing evaluations focus on tree reconstruction Ribeiro et al. (2023) and QA performance, avoiding considerations of whether the quality of model-generated trees aligns with the accuracy of the answers. This is largely because reasoning quality is difficult to assess. Evidence suggests humans think of deductive validity less stringently than formal symbolic axioms; they accept explanations that are more or less incomplete Sulik et al. (2021); Tan (2022). Attempts to annotate datasets Tafjord et al. (2022); Clark et al. (2023) consequently suffer from inconsistent labels, as opinions differ between annotators over what constitutes validity. Without careful instruction, it is hard for annotators to determine where to set their threshold for acceptability on the spectrum from “strict deductive inferences” to “fallible common-sense inferences” Gubelmann et al. (2023). Thus, asking for a single binary decompositional entailment judgment yields a difficult and noisy annotation task (Figure 1, (A)). This leads us to pursue a more consistent protocol for measuring the goodness of one decomposition over another.
@@
Existing discourse surrounding NLI (e.g. Manning (2006)) has not touched on the sort of decompositional entailments encountered in a recursive tree search algorithm. Towards addressing this gap, we propose an evaluation centered on the notion that a valid decomposition is a valid argument for why the hypothesis should be believed. We take inspiration from the “Relevance, Acceptability, and Sufficiency” criteria for a logically good argument developed in the field of informal logic Johnson and Blair (1977); Groarke (2022), and design an approach to assessing entailment in a principled manner. We annotate items on ordinal scales targeting different facets of argument analysis, yielding a well-defined and consistent process. We use this protocol to collect over 1000 expert annotations. Through experiments, we find that models trained on previous compositional entailment datasets and LLMs like GPT fall short of human performance on the dataset, which we term RDTE (Recognizing Decompositional Textual Entailment).
@@
We also find that our annotation protocol can serve as a powerful LM prompting mechanism: Given the same instructions as our human annotators, GPT-4 OpenAI (2023) can serve as a “teacher” in a knowledge distillation pipeline that annotates reasoning traces extracted from a tree search algorithm. We release a large artifact (24K items per domain) of GPT-4’s annotations for this purpose. We illustrate the effectiveness of student models trained via this pipeline as the linchpin of TreeWise, a novel entailment tree engine that performs hypothesis proving via decompositional entailment search. Inspired by the recent Nellie engine Weir et al. (2023), TreeWise leverages in-context learning, forward chaining, branch consolidation, and most importantly, improved decompositional entailment recognition. It surpasses tree-producing approaches on established benchmarks like EntailmentBankQA, but also adapts to complex tasks such as HotpotQA that require reasoning over less structured knowledge sources like Wikipedia. Using the knowledge-distilled models as reasoning verifiers improves TreeWise’s QA accuracy and raises the quality of the trees it produces. Our contributions are thus:
@@
  • A new entailment challenge set, RDTE, crafted via an informal logic-inspired protocol, tasking models to validate hypothesis decompositions
@@ 2 Decompositional RTE
2.1 Task Definition
An entailment tree comprises one or multiple hypothesis decompositions. Each constitutes one instance of decompositional RTE, which we define as follows: Given the question $Q$, hypothesis $h$, and premises $[p_{1},\dots,p_{n}]$, would proving the premises be contextually sufficient to prove $h$?
As described below, we collect different ordinal scores for each item using criteria from informal logic. Each facet informs the ultimate sufficiency label, for which we evaluate models in §4.
@@ 2.2 Background: RAS Criteria
A seminal framework developed by informal logicians to replace strict deductive logic criteria is known as RAS: Relevance, Acceptability, and Sufficiency Johnson and Blair (1977). Much like RTE work, RAS assesses whether a hypothesis can be accepted on the basis of a list of premises. Noting that each element is subject to substantial academic debate (different systems of informal logic define them differently), we review each criterion as we interpret them for this work:
Relevance of premise A concerning conclusion B is defined as whether the truth of A makes a difference to the truth of B. Blair (2012). The extent of this difference can vary; e.g. whether (A) “the earth has oxygen” is true technically has relevance to (B) birds can fly (since birds breathe oxygen), but is less relevant than (A’) birds have wings. Relevance is of particular interest to a recursive reasoner, as one would not want to waste time proving an irrelevant statement like (A) when proving (B).
@@
Sufficiency is “the property of an argument’s premises of supplying all the grounds that are needed to make it reasonable to believe its conclusion.” Johnson and Blair (1977). This is left intentionally vague; Blair (2012) notes “the criterion of sufficiency, for justificatory arguments, is best seen as a placeholder for whatever version and standards of sufficiency are appropriate for the particular situation in question.” In this way, we should consider sufficiency to be a question and problem-specific criterion: the grounds for valid entailment are inherently domain-specific. Blair (2012) notes the dependent relationship between sufficiency and the other two criteria: sufficient presumes acceptability and relevance.
@@ 2.3 Implementing RAS for RTE Annotation
We draw inspiration from these criteria to construct a protocol that investigates a decomposition like the ones shown in Figure 1 for each RAS component in turn. An important aspect of these normative terms is that they are all scalar: a premise can be more or less relevant and acceptable, and the sufficiency of an argument composed of premises could always be strengthened by adding more premises. In a departure from works such as Tafjord et al. (2022) who collect binary factuality and “reasoning correctness” judgments, we collect RAS judgments on an ordinal scale Zhang et al. (2017) from 1 to 5, where each score is assigned a specific set of conditions. We also collect an ordinal judgment for redundancy, which is not strictly a component of RAS assessment; we observe that redundant premises are particularly problematic for entailment tree search, as proving the same information twice wastes search budget. We implement redundancy as “conditional irrelevance”:
@@
The sufficiency label most directly resembles that of a typical RTE label, and is what we evaluate models on in later sections. Following Blair (2012)’s observation, the sufficiency judgment is higly dependent on the other criteria; our annotation directions, found in §B, include a substantial list of 30+ conditions around a decomposition that indicate what the sufficiency score should be based on other criteria scores, e.g. “Did you give any fact a 3/5 or lower for redundancy, factuality, or relevance? (Yes = max 4 sufficiency)”.
@@ 3 Data Collection
We seek to construct a dataset of decompositional entailment judgments that are of the kind that a reasoning system might come across while performing QA. Following Tafjord et al. (2022), we found it apt to annotate decompositions generated during a model’s reasoning search traces over training questions. We use the backward chaining process of the tree search algorithm TreeWise, which is introduced in §5. Each item is a hypothesis and a set of 2-3 premises that the model proposes might conjunctively entail the hypothesis in the context of a given question. To collect a representative sample of hypotheses, we pull from three classes:
@@
    Top-level correct hypotheses representing the right answer options for multiple-choice questions. These are important to annotate so that models are right for the right reasons.
@@
    Recursive correct hypotheses LLM-generated to decompose the top-level correct hypotheses
@@
    Top-level incorrect hypotheses representing the incorrect answer deemed closest to correct by GPT-4. These are important to annotate so that models are not wrong for the wrong reasons.
@@
3.1 Annotation Process
Four of this paper’s authors, all highly proficient or native English speakers, annotated 1000 decompositions total with two-way redundancy. One author annotated all items. We first annotated several examples together to develop a list of 30+ condition checks that indicate certain RAS scores. We then annotated the rest of the dataset independently.
While the dataset is annotated for sufficiency on a 5-point scale, we found a threshold of $\geq 4$ fitting to evaluate models under binary entailment metrics. This corresponds to items for which the only acceptable flaws are (A) a 3rd redundant premise in an otherwise perfect 2-premise entailment or (B) minor missing information that does not affect lay reasoning. To arrive at a single clean entailment label for evaluation, we reconciled disagreements by discussing all items for which its two annotated sufficiency scores were on either side of 3.5. A vast majority disagreements were due to human error, e.g. missing a condition or not recognizing a particular flaw in reasoning.
Comparison to Existing Data Prior to reconciling disagreements, we measured raw annotation agreement and compared it to recent attempts to collect decompositional entailment labels. Tafjord et al. (2022) collected annotations for 3.7K decompositions generated by Entailer, while Clark et al. (2023) did so for 24.4K items generated by GPT3.444The BaRDa dataset comprises a specifically filtered subset of these two datasets that exhibit maximal annotator agreement. We obtained the pre-filtered data from the authors for the purposes of comparison with our work. The instructions given to annotators for these datasets are very high-level, simply asking whether the “reasoning goes wrong.” We believe this creates a much lower threshold for acceptability than the RDTE protocol; these datasets are majority labeled “entailment,” but on manual inspection we found numerous items that exhibited redundant, irrelevant, and fallacious arguments.
Nevertheless, RDTE has a higher internal annotator consistency rate than these datasets, with a raw agreement rate 9% higher and Krippendorf $\alpha$ 19 points higher than the previous best.
RDTE (ours) Entailer GPT3
Raw agreement .79 .70 .61
Krippendorf $\alpha$53 .34 .31
@@ 3.1 Annotation Process
P1: Frank Vincent appeared in The 2005 film Remedy. 5 5
P2: Frank Vincent appeared in The Sopranos and several mob movies 5 5
P3: Ralph De Vito directed several mob movies. 5 5
@@
P1: Changes in the Grand Canyon’s landscape include becoming wider. 5 5 5
P2: Erosion is one of the processes that can change a landscape. 5 5 5
P3: The Grand Canyon is a landscape. 5 3 5
@@
P1: Ordos City is predominantly rural. - 3 5
P2: Predominantly rural areas in China are often found in the western part of the country. - 3 5
@@
Figure 3: Example RDTE annotations.
3.2 RDTE Analysis
@@ 4 RDTE Evaluation
We elicit sufficiency judgments over the 1000 gold-annotated RDTE items from a series of methods based on the RDTE protocol as well as existing methods for judging decompositional entailment. We test them on the gold sufficiency scores converted to binary labels by assiging 4 and 5s to positive entailment. Factuality, relevance, and redundancy predictions were not directly evaluated.
GPT Methods We test the following prompt-based methods, using both GPT-4 and ChatGPT.555gpt-4-0613 and gpt-3.5-turbo-1106 All prompts can be found in the appendix.
  • An ICL prompt containing the RDTE annotation rubric and 4 example batches (10-15 decompositions per hypothesis) and a Zero-Shot prompt containing only the RDTE rubric.
  • The prompts used by Clark et al. (2023) to evaluate models on the BaRDa dataset. These are separate prompts for the entailment and premise factuality judgments (see Figure 10).
Existing Methods We test various RTE models from recent entailment tree-based reasoners: the T5-based Entailer-11B Tafjord et al. (2022) and the similar Nellie-3B model Weir et al. (2023), as well as the fine-tuned RoBERTa classifier used by the Nellie system as an entailment filter. We also take an off-the-shelf (OTS) RoBERTa finetuned for NLI using a suite of 600 datasets666huggingface.co/sileod/deberta-v3-large-tasksource-nli that includes QASC Khot et al. (2020) and ARC.
Knowledge Distillation on Silver Data Search algorithms like TreeWise make hundreds of calls to NLI models for every question, making large, slow models like GPT-4 unrealistic for use as an entailment filter. We instead explore using knowledge distillation to train smaller student models to imitate GPT-4’s performance. We extracted 24K RDTE judgments from GPT-4 over 2.3K hypotheses in each of the ARC and Hotpot domains. We release these large silver datasets for future work. [^7] GPT-4 obtains a .61 and .44 Spearman $\rho$ with humans on the ARC and HotpotQA RDTE splits, respectively.
We used the silver data to fine-tune (A) the RoBERTa model used in the Nellie engine, and (B) ChatGPT using OpenAI’s fine-tuning API. We trained the former to predict the binarized silver sufficiency judgments and the latter to replicate GPT-4’s exact answers to the RDTE prompt. For token efficiency, the GPT-4 teacher and ChatGPT student perform classification over a batch of decompositions for a single hypothesis in each prompt. [^8] Batching barely affects performance on RDTE (Table 3).
@@
ChatGPT (RDTE Zero-Shot) 40 38
ChatGPT (BaRDa) 43 34
@@
ChatGPT 48 (+5) 51(+13)
RoBERTa 66 (+21) 56(+12)
Table 1: RDTE entailment results (F0.5 score). We find that RDTE prompting of GPT-4 greatly outperforms existing approaches to compositional entailment, including BaRDa prompting and fine-tuned approaches. Knowledge disillation from GPT-4 (RDTE Zero-Shot) improves student models by 5-21 points, at times outperforming GPT-4 itself. $\dagger$/$\ddagger$ / ‡ denote the student models and the best performing analogous non-distilled models.